OpenStudy (anonymous):

Find an equation for the nth term of the arithmetic sequence. -17, -13, -9, -5, ...

OpenStudy (nincompoop):

do you know the formula?

OpenStudy (john_es):

You must find the distance first, \[-13-(-17)=4\]And apply the usual arithmetic formula, \[a_n=a_0+d(n-1)=-17+4(n-1)\]

OpenStudy (anonymous):

I know, so far this is what I've obtained: an = -17 + 4

OpenStudy (anonymous):

Wait, is it an = -17 + 4(n - 1)

OpenStudy (john_es):

Yes :)

OpenStudy (nincompoop):

something like that